# captionUnknown document class (or package)

### v1

@@ 5.3 End-to-End Architectures for 3D Vision & Language
Transformer models pre-trained on large 3D-text datasets learn powerful joint representations that bridge the visual and language modalities. 3D-VisTA [343] is a transformer model that employs self-attention to jointly model 3D visual and text data, enabling effective pre-training on objectives like masked language/object modeling and scene-text matching. UniT3D [227] takes a unified transformer approach, combining a PointGroup 3D detection backbone, BERT text encoder, and multi-modal fusion module, with joint pre-training on synthetically generated 3D-language data. SpatialVLM [344] takes a different tack, and co-trains VLMs on a large synthetic 3D spatial reasoning dataset, boosting performance on 3D spatial visual question answering tasks and enabling applications like chain-of-thought reasoning for robotics. Multi-CLIP [345] pre-trains a 3D scene encoder to align scene features with CLIP’s text and image embeddings, aiming to transfer CLIP’s knowledge for improved 3D understanding on tasks like visual question answering.
In addition to pre-training approaches, researchers have explored architectures that unify 3D perception with language capabilities in an end-to-end framework. D3Net [27] combines dense captioning and visual grounding with a 3D object detector, a speaker for generating captions from detections, and a listener for discriminating objects using the captions. Uni3DL [28] operates on point clouds, with modules for text encoding, point encoding, semantic/mask prediction, and diverse task outputs like segmentation, detection, grounding, and captioning. InstanceRefer [346] uses panoptic segmentation and linguistic cues to filter instance candidates based on language descriptions for visual grounding tasks in 3D point clouds, while LanguageRefer [347] combines language embeddings with spatial embeddings from 3D bounding boxes. 3DVG-Transformer [348] also tackles 3D grounding in point clouds, with a coordinate-guided contextual aggregation module and multiplex attention for effective feature fusion.
Dataset     Object Captioning     Scene Captioning     Dense Captioning     Single Object Grounding     Multiple Object Grounding     Question Answering     Situated Question Answering     Dialogue     Task Planning     Real     Synthetic     Object     Indoor     Outdoor     Human-Annotated     Model-Annotated     Template-Based
Cap3D [349]
Text2Shape [350]
SceneVerse [351]
nu-Caption [271]
nu-Grounding [271]
ScanRefer [218]
ReferIt3D [228]
Multi3DRefer [230]
Chat-3D v2 [172]
EmbodiedScan [352]
ScanEnts3D [353]
WildRefer [354]
RIORefer [355]
ARKitSceneRefer [356]
ScanERU [229]
DenseGrounding [231]
ScanQA (Azuma et al.[234]
ScanQA (Ye et al.[357]
3DMV-VQA [236]
NuScenes-QA [235]
CLEVR3D [358]
SQA-3D [233]
3D-LLM [153]
ScanScribe [294]
M3DBench [359]
GPT4Point [268]
LAMM [360]
@@ 6 Datasets
We now provide a high-level overview of the datasets that are used to train and evaluate 3D vision-language models. In Table III, we provide a list of datasets alongside the tasks they are used for, as well as information about the 3D scans and annotations. In Fig. 4, we present these datasets on a timeline, showing where each dataset sourced 3D information from. Current 3D vision-language datasets are almost exclusively generated by taking existing 3D vision datasets, and applying human, model or templated annotations to samples. As seen in Table III, a majority of existing datasets focus on real, indoor scenes, and this is partially explained by observing that the majority of existing datasets use 3D scans from ScanNet [361] and 3RScan [295]. Many of the datasets presented here share the same 3D data, and instead differ primarily through their choice of annotation strategy, and the 3D vision-language task they are designed to be used for.
@@
Datasets for 3D navigation and manipulation using language are often designed around specific requirements and have a large overlap with existing bodies of research. We refer readers to existing survey papers [362, 363] for an overview of these datasets. Similarly, for text-to-3D generation datasets, we direct readers to the recent survey by Lee et al. [243]. We omit further discussion here due to prior extensive coverage, and because many methods use 2D vision-language data rather than 3D-specific datasets.
@@
Figure 4: Timeline of datasets. A timeline showing how existing datasets are combined and annotated to form new datasets for 3D vision language tasks. Datasets in orange are foundational 3D datasets without language annotations and datasets in blue are the annotated datasets used in 3D vision language tasks. Many existing datasets use 3D data from the same sources, such as ScanNet and 3RScan, and instead differ primarily through their choice of annotation strategy and targeted 3D vision task. Note that WildRefer also introduces new 3D data and annotations for vision-language tasks.
Cap3D [349] is a 3D object captioning dataset developed upon 660k objects from the Objaverse [293] dataset. It is constructed by generating 2D image captions from multiple views of the 3D object, and consolidating those captions using image-text alignment and LLMs.
Text2Shape [350] is a human-captioned form of the 8,447 tables and 6,591 chairs found in ShapeNet [296], combined with a procedurally generated dataset of primitive shapes labeled with template-based captions. It was originally used for generative text-to-3D shape tasks.
SceneVerse [351] is a large-scale, multi-purpose dataset of annotated scenes made by compiling 68k scenes from existing 3D datasets. SceneVerse contains a total of 2.5 million vision-language pairs for object captioning, scene captioning and generating relative descriptions, primarily generated through the use of 3D scene graphs and LLMs.
nu-Caption [271] is a captioned version of 420k LiDAR scans from the nuScenes [364] dataset, annotated using GPT-4 and 2D MLLMs. Captions include general scene descriptions, detailed descriptions of objects and their relationships, and the identification of potential risks on the road.
nu-Grounding [271] builds on nu-Caption by focusing on the task of grounding, using the annotations from nuScenes to to create 280k pairs of questions and answers for visual grounding and grounded captioning.
ScanRefer [218] introduces the task of 3D RGB-D grounding using a natural language expression by creating 51,583 human-annotated ‘referring expressions’ of 11,046 objects found in 800 scenes of ScanNet [361]. The input consists of a point cloud of a scanned 3D scene alongside a free-form description of a specified target object, and the output is the corresponding bounding box of the object. ScanRefer provides an evaluation server and online benchmark  [^2] https://kaldir.vc.in.tum.de/scanrefer_benchmark/ to enable comparison between different methods.
ReferIt3D [228] introduces a number of datasets (Nr3D, Sr3D and Sr3D+) consisting of objects from 707 ScanNet scenes. Similarly to ScanRefer, these object have been annotated with referring expressions, with a focus on queries where the scene contains multiple instances of the target class, and the referring expression is required to disambiguate between them. Nr3D contains 41,503 human-annotated, free-form utterances to refer to objects in 3D scenes, Sr3D contains 83,572 template-based utterances, and SR3D+ is a version of Sr3D with augmented utterances. ReferIt3D also provides an evaluation server and online benchmark  [^3] https://referit3d.github.io/benchmarks.html to enable comparison between different methods.
Multi3DRefer [230] is a modified version of the ScanRefer dataset. Instead of referring expressions that always refer to one object in the scene, Multi3DRefer contains 6688 zero-target, 42,060 single target and 13178 multiple target referring descriptions collected for 11,609 objects in 800 ScanNet scenes. ChatGPT [286] is also used to rephrase the referring expressions.
Chat-3D v2 [172] is another modified form of ScanRefer, where the referring expressions from 705 scenes in ScanNet are used to construct scene captions that describe the relationships between objects in the scene. The scene captions are generated using GPT-4 [365] by providing the model with ground-truth information about objects. The generated captions contain references to explicit ‘object identifiers’ that directly represent each object in the scene.
EmbodiedScan [352] is an annotated combination of Matterport3D [285], 3RScan [295] and ScanNet [361], designed as a multi-modal, ego-centric dataset for 3D scene understanding. Segment Anything [366] and other annotation tools are used to provide 3D bounding boxes, semantic occupancy, and 970k template-based language descriptions across a total of 5185 scenes.
ScanEnts3D [353] extends both ScanRefer [218] and ReferIt3D [228] by using professional annotators to link every object mentioned in referential sentences to their respective instances within a 3D scene. In the original paper this dataset is only used for training purposes, where it is found to improve the model’s performance on other visual grounding and captioning datasets.
WildRefer [354] presents the STRefer and LifeRefer datasets, emphasizing human-centric in the wild settings with comprehensive 3D and linguistic human-annotations for 3D grounding. STRefer contains 5,458 referring expressions for objects in 662 scenes from the STCrowd [367] dataset, whereas LifeRefer contains 25,380 referring expressions for objects in 3,172 scenes from a new set of 3D scans acquired for this dataset.
RIORefer [355] is a human-annotated version of 3RScan [295], used for 3D grounding. It consists of 63k descriptions of objects across 1,380 scenes. This dataset is introduced as a way of testing the cross-dataset generalization abilities of models, such as in the proposed “ScanRefer to RIORefer generalization” and “RIORefer to ScanRefer generalization” tasks.
ARKitSceneRefer [356] is an annotated version of ARKitScenes [368] emphasizing the 3D grounding of small, everyday objects found in real-world indoor environments. It contains 15k descriptions of objects found in 1,605 scenes.
ScanERU [229] is a modified and human-annotated version of ScanNet that combines 46k referring expressions from ScanRefer with 706 ScanNet scenes that have been modified to include a 3D human model gesturing to the referred object, using locations designated by human annotators.
DenseGrounding [231], like Multi3DRefer, seeks to extend the task of 3D grounding to include multiple objects, however instead of a single referring expression referring to multiple objects, each input is a combined paragraph of referring queries which each correspond to a single object. These paragraphs are constructed using the nearest neighbors of random objects from ScanRefer and ReferIt3D, and combining their referring expressions together to form a paragraph.
ScanQA (Azuma et al.) [234] is an annotated version of ScanNet with 41k question-answer pairs generated across 800 scenes. Questions are automatically generated using the referring expressions in ScanRefer and then refined by human annotators, whereas answers are derived entirely from human-annotators. This is the dataset that is typically referred to by the name “ScanQA”.
ScanQA (Ye et al.) [357] was released contemporaneously with ScanQA (Azuma et al.) and is also a human-annotated form of ScanNet to be used as a 3D question answering dataset. Ye et al.contains 10k question-answer pairs for 806 ScanNet scenes. While Azuma et al.initially generates questions using referring expressions from ScanRefer, Ye et al.use human annotators for the entirety of question creation.
3DMV-VQA [236] is an annotated version of 5k scenes from the Habitat-Matterport 3D Dataset (HM3D) [369], using the semantic information from HM3DSem [370] to generate 50k questions of four types: “concept”, “counting”, “relation” and “comparison”. These questions are generated as templates, and then transformed into natural language questions.
NuScenes-QA [235] consists of 34k scenes from nuScenes [364] annotated with 460k template-style question-answer pairs generated using a constructed scene graph. Questions are divided into 5 types: “existence”, “counting”, “query-object”, “query-status” and “comparison”, and can include spatial reasoning.
CLEVR3D [358] is an annotated form of 3RScan [295] designed for indoor 3D QA tasks. Template-based questions and answers are generated using the scene graph annotations from 3DSSG [371]. Initially, 44k questions are generated for 1,333 scenes, however they use a “compositional scene manipulation” technique to randomly replace objects in the scene graphs with objects from a generated object pool, artificially leading to 171k questions across 8,771 scenes.
SQA-3D [233] is another human-annotated version of ScanNet, proposing the problem of “situated question answering”, where each query consists of a description of the agent’s current location and situation, and a query relating to the scene. SQA-3D contains 33.4k questions across 650 scenes.
3D-LLM [153] is an annotated version of Objaverse, ScanNet and HM3D/HM3DSem designed to handle a broad variety of tasks involving 3D vision and language. BLIP-2 [195] and ChatGPT [286] are used to generate 300k instances of ‘3D-language data’ that are used across a range of 3D vision-language tasks.
ScanScribe [294] is an annotated version 1,185 scenes from ScanNet and 3RScan designed to act as a large-scale 3D scene-text pair dataset for pre-training 3D vision-language models. It utilizes ScanQA, ScanRefer and ReferIt3D for ScanNet data, and uses GPT-3 prompting to generate annotations for 3RScan to create a total of 278k scene descriptions. To increase the diversity of the dataset, scenes are synthetically generated by random replacing 10% of the objects in scenes with a same-category object from Objaverse.
M3DBench [359] is a multi-modal instruction-following dataset consisting of 327k instruction-response pairs across a broad range of tasks. Data is collected from a large set of existing datasets (see Fig. 4), alongside using GPT-prompting to generate additional annotations.
GPT4Point [268] is an annotated form of over 1 million objects from Objaverse-XL, designed for 3D Captioning, 3D QA and other 3D tasks. Annotations are generated automatically through a hierarchical pipeline that fuses and improves captions from multiple views.
LAMM [360] is a large scale multi-modal instruction-tuning dataset that covers tasks for both 2D and 3D vision. The 3D Data is sourced from 3RScan [295], CLEVR3D [358], 3DSSG [371] and ShapeNet [296], and is annotated using GPT [365] and template-based responses.
@@ 7 Challenges and Opportunities
Despite progress in integrating LLMs with 3D data, challenges remain in data representation, computational efficiency, and benchmarks, necessitating innovative solutions.
Representation Choice has a large impact on the performance of 3D vision language models. Currently, point clouds are predominantly used to represent both indoor (e.g. vertices of meshes) and outdoor (e.g. LiDAR point clouds) environments due to their simplicity and neural network compatibility. However, they struggle to capture the fine details crucial for accurate, rich spatial models. Developing novel 3D scene representations that more effectively bridge the gap between spatial information and language could unlock new levels of understanding and interaction. By finding innovative ways to encode language and semantic information in 3D representations, such as using distilled language and semantic embeddings [30, 316], could help bridge the gap between these two modalities.
Computational Demands of both 3D data processing and LLMs pose significant challenges. Scalability remains a concern as the complexity of 3D environments and the size of language models increase. Advances in LLM architectures designed for adaptivity and computational efficiency could significantly broaden their application scope.
Improving Benchmarks is essential for comprehensively evaluating and advancing the capabilities of multi-modal LLMs in 3D tasks. Current benchmarks have limited scope, especially regarding 3D reasoning, hampering evaluation of spatial reasoning skills as well as development of 3D decision-making/interaction systems. Moreover, the metrics in use today fall short of capturing the full spectrum of LLMs’ capabilities in 3D environments. Crafting task-specific metrics that more precisely measure performance across diverse 3D tasks is essential. Finally, the granularity of current scene understanding benchmarks is overly simple, limiting insight into complex 3D environment comprehension. A more diverse array of tasks is required.
Safety and Ethical Implications are imperative to consider when employing LLMs for 3D understanding. LLMs can hallucinate and output inaccurate, unsafe information leading to erroneous decisions in critical 3D applications. Moreover, LLMs often fail in ways that are unpredictable and challenging to interpret. They may also inherit societal biases present in training data, disproportionately disadvantaging certain groups when making predictions in real-world 3D scenes. It is crucial to approach the use of LLMs in 3D contexts with caution, employing strategies to create more inclusive datasets, robust evaluation frameworks for bias detection and correction, and mechanisms to minimize hallucinations, ensuring responsible and equitable outcomes.

### v2

@@ 5.3 End-to-End Architectures for 3D Vision & Language
Transformer-based 3D-text models learn joint visual-language representations. 3D-VisTA [260] models both modalities with self-attention and is pre-trained on masked modeling and scene-text matching. UniT3D [110] combines PointGroup, BERT, and multi-modal fusion, trained on synthetic 3D-text data. SpatialVLM [261] trains VLMs on synthetic 3D spatial reasoning tasks for improved 3D spatial VQA and chain-of-thought reasoning. Multi-CLIP [262] aligns 3D scene encodings with CLIP’s embedding space for better cross-modal transfer. Lexicon3D [263] benchmarks a wide range of foundation models across 3D vision-language tasks, identifying key strengths (e.g., DINOv2 for vision, diffusion for grounding) and challenges in multimodal alignment.
End-to-end 3D-language models unify perception, grounding, and captioning. D3Net [264] includes a detector, a speaker module for captioning, and a listener for grounding. Uni3DL [265] operates directly on point clouds with shared encoders for text and points, predicting outputs for segmentation, detection, grounding, and captioning. InstanceRefer [266] and LanguageRefer [267] combine linguistic and spatial cues for grounding in 3D point clouds. 3DVG-Transformer [268] uses coordinate-guided attention and contextual feature fusion for grounding descriptions to 3D objects. SAGA [269] enables prompt-based segmentation using 3D Gaussian splats with SAM-derived affinity fields, supporting real-time open-vocabulary interaction across granularities.
6 Datasets and Benchmarking
Dataset Object Captioning Scene Captioning Dense Captioning Single Object Grounding Multiple Object Grounding Question Answering Situated Question Answering Dialogue Task Planning Real Synthetic Object Indoor Outdoor Human-Annotated Model-Annotated Template-Based # Language Pairs # Scenes
Cap3D [270] 660k 660k
Text2Shape [271] 75k 15k
SceneVerse [272] 2.5M 68k
nu-Caption [155] 420k -
nu-Grounding [155] 280k -
ScanRefer [101] 51k 800
ReferIt3D [111] 125k 707
Multi3DRefer [113] 60k 800
Chat-3D v2 [81] 705 705
EmbodiedScan [273] 970k 3.4k
ScanEnts3D [274] 84k 700
WildRefer [275] 30.8k 3.8k
RIORefer [276] 1.4k 63k
ARKitSceneRefer [277] 15k 1.6k
ScanERU [112] 46k 706
PhraseRefer [278] 170k 700
DenseGrounding [114] - -
3DVQA [279] 484k 707
ScanQA (Azuma et al.) [58] 41k 800
ScanQA (Ye et al.) [280] 10k 806
3DMV-VQA [118] 50k 5k
NuScenes-QA [117] 460k 34k
CLEVR3D [281] 171k 8.7k
SQA-3D [59] 33.4k 650
MSQA/MSNN [282] 251k 1.7k
3D-LLM [115] 200k 900
ScanScribe [202] 278k 1.2k
M3DBench [283] 327k -
GPT4Point [154] 1M 1M
LAMM [284] - -
3D-GRAND/3D-POPE [285] 6.2M 40k
LEO [152] - -
3DMIT [157] 75k 1.5k
Spartun3D [193] 133k -
3D-CoT [286] 1.51M -
Beacon3D [287] 5k 30
SPAR-Bench [288] 7M 4k
SpatialVQA [172] 1.3k -
Space3D-Bench [289] 1k -
City-3DQA [290] 450k -
MMScan [291] 6.9M 5.2k
Reason3D 2.5k -
LEO-VL 700k -
SpaCE-10 6k 811
@@ 6 Datasets and Benchmarking
We now provide a high-level overview of the datasets that are used to train and evaluate 3D vision-language models. In Table III, we provide a list of datasets alongside the tasks they are used for, as well as information about their 3D data and annotations. In Fig. 5, we present these datasets on a timeline, showing where each dataset sourced 3D information from. Current 3D vision-language datasets are almost exclusively generated by taking existing 3D vision datasets, and applying human, model or templated annotations to samples. As seen in Table III, a majority of existing datasets focus on real, indoor scenes due to using 3D scans from either ScanNet [57] or 3RScan [203]. Many of the datasets presented here share the same 3D data, and instead differ primarily through their choice of annotation strategy, and the 3D vision-language task they are designed to be used for. Among outdoor datasets, the majority use nuScenes [292], which focuses on LiDAR data for autonomous driving.
There are three current strategies for collecting language data for these datasets: human-based, template-based and model-based annotation. Prior to ChatGPT [191], language data was collected from human annotators labelling data with natural language, or by having humans construct natural language ‘templates’, where appropriate classes and spatial references are taken from the dataset (such as from scene graphs) and placed into the template phrases. However, human-based annotation is expensive and time-consuming, and template-based annotation can have limited linguistic diversity. Recently, datasets have been using LLMs to analyze and annotate 3D data, either by asking the LLM to analyze a generated text description of each 3D scene, or by using 2D vision or vision-language models to analyze multiple 2D images of the scene and collating the responses. This has allowed 3D vision language datasets to contain orders of magnitude more language annotations, but can fail when the LLM does not generate accurate and naturalistic annotations for the samples. See Table III for a breakdown of how each dataset is annotated.
Due to LLMs’ ability to act as multi-task learners, several recent datasets [154, 157, 152, 202, 283, 284, 285] have focused on collecting data for many different tasks, attempting to unify the input and output formats for each task (see section 4.3). These datasets typically draw 3D data from a greater range of datasets to expand data size and diversity (see Fig. 5). There are a few strategies to improve the data diversity by leveraging multiple datasets, such as randomly replacing 10% of the objects in scenes with a same-category object from Objaverse [202], or randomly replacing objects in the scene graphs with objects from a generated object pool [281]. Many datasets can also be used beyond their intended task. Dense captioning methods such as Scan2Cap [100] can use the referring expressions from grounding datasets like ScanRefer [101] to create the captions needed to train a model. Datasets that were designed for text-to-3D generation tasks, such as [271], can also be used for 3D-to-text captioning tasks. Other datasets are introduced as a way of testing the cross-dataset generalization abilities of models [276].
@@
Datasets for 3D navigation and manipulation using language are often designed around specific requirements and have a large overlap with existing bodies of research. We refer readers to existing survey papers [293, 294] for an overview of these datasets. Similarly, for text-to-3D generation datasets, we direct readers to the recent survey by Lee et al. [124]. We omit further discussion here due to prior extensive coverage, and because many methods use 2D vision-language data rather than 3D-specific datasets. In Fig. IV, we provide a quantitative comparison between a number of existing 3D-VLM methods on the common testing datasets ScanRefer [101], Multi3DRefer [113], Scan2Cap [100], ScanQA [58] and SQA3D [59]. We include these results to highlight the variety of metrics and datasets used even among similar methods.
@@
Figure 5: Timeline of datasets. A timeline showing how existing datasets are combined and annotated to form new datasets for 3D vision language tasks. Datasets in orange are foundational 3D datasets without language annotations and datasets in blue are the annotated datasets used in 3D vision language tasks. Note that WildRefer also introduces new 3D data and annotations for vision-language tasks.
Model ScanRefer Multi3DRefer Scan2Cap ScanQA (val) SQA3D (test)
Acc@0.25 Acc@0.5 F1@0.25 F1@0.5 B-4@0.5 C@0.5 M@0.5 R@0.5 M R C EM B-1 B-4 EM EM-R
Specialists
ScanRefer [101] 37.3 24.3 - - - - - - - - - - - - - -
Scan2Cap [100] - - - - 22.4 35.2 22.0 44.8 - - - - - - - -
Vote2Cap-DETR++ [109] - - - - 34.5 61.8 26.2 54.4 - - - - - - - -
ScanQA [58] - - - - - - - - 13.1 33.3 64.9 21.1 30.2 10.1 47.2 -
SQA3D [59] - - - - - - - - 13.5 34.5 - - 30.5 11.2 46.6 -
D3Net [264] - 37.9 - 32.2 35.7 62.6 25.7 53.9 - - - - - - - -
3D-ViSTA [202] 50.6 45.8 - - 34.0 66.9 - - 13.9 35.7 69.6 22.4 - - 48.5 -
3DVG-Transformer [268] 45.9 34.5 - 25.5 - - - - - - - - - - - -
SIG3D [151] - - - - - - - - - - - - - - 52.6 -
Generalists (Point-Based Methods)
Chat-3D [80] - - - - - - - - - - 53.2 - - - - -
Chat-3D v2 [81] 42.5 38.4 45.1 41.6 31.8 63.9 - - 16.1 40.1 87.6 - 38.4 7.3 54.7 -
LL3DA [19] - - - - 36.0 62.9 26.0 55.0 15.9 37.3 76.8 - - 13.5 - -
Scene-LLM [147] - - - - - - - - 16.6 40.0 80.0 27.2 - - 53.6 -
LEO [152] - - - - 38.2 72.4 27.9 58.1 16.2 39.3 101.4 21.5 - 11.5 50.0 52.4
Chat-Scene [156] 55.5 50.2 57.1 52.4 36.3 77.1 - - 18.0 41.6 87.7 21.6 43.2 14.3 54.6 57.5
3D-LLaVA [171] 51.2 40.6 - - 36.9 78.8 27.1 57.7 18.4 43.1 92.6 27.0 - 17.1 54.5 56.6
Robin3D [163] 60.8 55.1 64.9 59.7 38.4 87.2 - - 19.2 44.0 - - - - 56.9 59.8
PerLA [164] - - - - - - - - 17.4 39.6 78.1 - - 14.5 - -
Grounded-3DLLM [295] 47.9 44.1 45.2 40.6 35.5 70.6 - - 15.2 37.1 72.7 - - - -
Inst3D-LLM [168] 57.8 51.6 58.3 53.5 38.3 79.7 - - - - 88.6 24.6 - - - -
Spartun3D-LLM [193] - - - - - - - - - - - - - - 55.0 -
LSceneLLM [167] - - - - - - - - 18.0 40.8 88.2 - - - - -
Generalists (Multi-View Image-Based Methods)
3D-LLM [115] 30.3 - - - - - - - 14.5 35.7 69.4 20.5 39.3 12.0 - -
LLaVA-3D [162] 54.1 42.4 - - 41.1 79.2 30.2 63.4 - - 91.7 27.0 - - 55.6 -
Video-3D LLM [165] 58.1 51.7 58.0 52.7 41.3 83.8 28.5 61.7 19.8 49.0 102.1 30.1 47.1 16.2 58.6 -
3DRS [176] 62.9 56.1 60.4 54.9 41.6 86.1 - - - - 104.8 30.3 - - 60.6 -
GPT4Scene-HDM [166] 62.6 57.0 64.5 59.8 - - - - 18.9 46.5 96.3 - 44.4 15.5 60.6 63.3
VeBrain [296] 66.4 60.2 67.8 62.7 - - - - 20.1 48.2 101.5 - 47.7 17.3 - -
VG LLM [173] 51.0 45.1 - - 38.7 74.1 28.0 61.0 - - - - - - - -
Ross3D [170] 61.1 54.4 59.6 54.3 43.4 81.3 30.3 66.9 20.9 50.7 107.0 30.8 - 17.9 63.0 65.7
LEO-VL [175] - - - - - - - - 17.4 39.6 78.1 - - 14.5 - -
Spatial-MLLM [172] - - - - - - - - 18.4 45.0 91.8 - 44.8 14.8 55.9 58.7
Human (from [297]) - - - - - - - - - - - 51.6 - - 90.1 -
Best Model 66.4 60.2 67.8 62.7 43.4 87.2 30.3 66.9 20.9 50.7 107.0 30.8 47.7 17.9 63.0 65.7
Model vs. Human Gap - - - - - - - - - - - 40.6% - - 30.1% -
TABLE IV: Quantitative comparison of different models on common 3D vision-language benchmarks. Results are from original papers where applicable. Best results in each category (Specialists and Generalists) are highlighted in bold. The bottom row shows the performance gap between the best AI models and human performance.
@@ 7 Challenges and Opportunities
Despite the progress made in integrating LLMs with 3D data, challenges remain in data representation, model training, evaluation and safety, necessitating further research.
7.1 Better 3D Representations for Language Alignment
A central challenge for 3D-LLMs is the choice of 3D data representation, as it directly affects both geometric fidelity and language alignment. Point clouds dominate due to their simplicity but lack fine-grained detail, while voxel grids and meshes provide richer geometry at the cost of high memory use or poor neural compatibility. Recent works have thus explored representations [231, 230, 298] that combine semantics with geometry, including differentiable formats such as NeRFs [49] and 3D Gaussian Splats [53], as well as VGGT-style geometry-aware encoders [200, 299]. Another promising direction is structured, part-level representations [300], which support hierarchical alignment between language and 3D across multiple granularities—from scenes, to objects, down to fine components.
7.2 Need for 3D-centric Pre-training and Architectures
Existing 3D-LLM methods typically rely on a frozen 3D encoder combined with an LLM that has been pretrained purely on text sequences. During fine-tuning, the majority of existing methods use alignment modules that transform native 3D features to sequences of 1D tokens that are directly aligned with the text embedding space of the pretrained LLMs. This reduction from inherently three-dimensional structures to one-dimensional text space leads to a substantial loss of spatial information, and these methods have little direct knowledge of 3D spatial structures beyond what can be inferred from linguistic priors. This highlights an important opportunity: to move toward 3D-centric pre-training and architectures that treat 3D structure as a first-class representation. Promising directions include introducing pretraining objectives that explicitly teach LLMs 3D relationships, such as masked autoencoding in a 3D token space, geometric consistency, and auxiliary reconstruction losses [170], and jointly pretraining the 3D encoder and the LLM so that both components co-adapt. In parallel, alignment modules should be redesigned from simple projection into language space toward bidirectional, structure-aware interfaces that better integrate 3D and text.
7.3 From Static to Dynamic Geometric Understanding
Most existing 3D-LLMs focus on static and coarse-grained comprehension of object categories, attributes (e.g. size, color), positions, and spatial relationships. However, many real-world tasks also require reasoning about dynamic properties from static 3D scans, such as predicting object affordances [301], functionality, and physical characteristics [302, 303] (e.g. material, fragility, deformability), as well as motion dynamics (e.g. free fall) and inter-object interactions (e.g. collision, friction [304]). Beyond this, fine-grained geometric reasoning, including symmetry, perspective, occlusion [137], and containment remain essential for precise spatial understanding. A next step is to move from inferring dynamics from static geometry to directly modeling dynamic 3D scenes (4D, i.e. 3D + time) [179, 305, 306], enabling LLMs to capture both spatial and temporal aspects of the physical world.
7.4 Acquisition of Higher Spatial Reasoning Ability
Recent progress such as OpenAI-o3 and DeepSeek-R1 [307] have shown that reinforcement learning (RL) can significantly enhance the reasoning ability of LLMs by providing reward signals beyond next-token prediction. However, current 3D-LLMs typically generate a single textual answer without engaging in step-by-step multi-hop reasoning or grounding their outputs to corresponding 3D structures. This lack of explicit reasoning chains and spatial grounding limits interpretability, which is essential for higher-order spatial reasoning. An important opportunity is to leverage RL for training 3D-LLMs with explicit reasoning traces and verifiable grounding. In particular, constructing 3D-specific chain-of-thought (3D-CoT) datasets could teach models to decompose problems into multi-step spatial reasoning processes. Together with well-designed reward functions and multi-stage training schemes (e.g. pre-training followed by RL-based post-training), 3D-LLMs may gradually acquire more robust, interpretable, and generalizable spatial reasoning skills, which is a largely unexplored direction.
7.5 Lack of 3D-text Paired Data as a Bottleneck
The success of LLMs is largely driven by massive text corpora, and 2D-VLMs rely on billions of image-text pairs such as LAION-5B [308]. In contrast, existing 3D vision-language datasets remain orders of magnitude smaller, with the largest containing only millions of 3D-text pairs (Tab. III). While VLMs can generate textual descriptions for 3D datasets, the scale of available 3D data is still far behind that of 2D images. Moreover, most existing annotations provide only coarse object-level descriptions, whereas fine-grained alignment across spatial, geometric, and semantic levels is critical for 3D-LLMs. This scarcity in both quantity and quality of 3D-text data is a major bottleneck for scaling and capability development in 3D-LLMs. Promising directions to alleviate this include: (1) using simulators [309] to synthesize diverse 3D data of objects, scenes, and interactions, combined with automated annotation tools for multi-granularity language alignment; and (2) leveraging 3D generative models [310, 311] to produce 3D content directly from textual descriptions, thereby expanding data scale while ensuring consistent 3D-text alignment.
7.6 Need for Accurate and Robust Evaluation Metrics
Current evaluations of 3D-LLMs largely follow a VQA-style protocol [58], where performance is measured only at the text level via similarity scores. However, complex 3D geometric relationships are often difficult to capture through language alone, and text-based answers do not necessarily reflect genuine spatial understanding. In particular, hallucinations pose a major issue: models may output the correct answer by exploiting linguistic priors rather than grounding their reasoning in the underlying 3D scene [312]. Existing metrics fail to detect this behavior, raising concerns about reliability. Going forward, the community needs more robust and interpretable evaluation methods that explicitly test 3D-grounded understanding and reasoning.
7.7 Safety Issues of 3D Embodied Agents
The ultimate goal of understanding 3D scenes is to enable actions, interactions, and transformations in the 3D world, where LLMs can play a crucial role. In Sec. 4.5, we discuss how current research leverages LLMs to empower 3D embodied agents in tasks such as planning, navigation, and manipulation, illustrating the feasibility of LLMs directly generating action outputs. However, a critical yet often overlooked challenge is the potential safety risks when LLMs generate physical actions. For instance, when controlling a robotic arm to assist an elderly person during feeding, the agent must carefully regulate its strength and speed to prevent harm. LLMs, however, may lack precise physical awareness or produce hallucinations, leading to excessive force or other hazardous actions. Thus, safety and reliability of 3D embodied agents must be ensured through more detailed mechanisms, including the introduction of more physical constraints [313], real-time feedback systems, and human-value alignment [314]. These measures are essential to achieving truly safe, reliable and trustworthy actions in complex 3D environments.